# GenoCraft: A Comprehensive, User-Friendly Web-Based Platform for High-Throughput Omics Data Analysis and Visualization

### v1

@@
[…]
${}^{1}$epartment of Pathology, Stanford University School of Medicine, Stanford, CA, 94305, USA
${}^{2}$ Bradley Department of Electrical and Computer Engineering, Virginia Tech, Blacksburg, VA, 24060, USA
${}^{3}$ Department of Computer Science, University of Southern California, Los Angeles, CA, 90089, USA
${}^{4}$ Computer Engineering, University of Illinois Urbana-Champaign, Urbana, IL, 61820, USA
${}^{5}$epartment of Computer Science and Engineering, University of Washington, 1410 NE Campus Parkway Seattle, WA 98195, USA
${}^{6}$epartment of Internal Medicine, Wake Forest University, Winston-Salem, NC, 27109, USA
${}^{7}$epartment of Computational Science and Engineering, Georgia Institute of Technology, Atlanta, GA, 30332, USA
$*{}^{*}$orrespondence: tfu42@gatech.edu, cvrechem@stanford.edu
‡Lead contact
@@
The surge in high-throughput omics data has reshaped the landscape of biological research, underlining the need for powerful, user-friendly data analysis and interpretation tools. This paper presents GenoCraft, a web-based comprehensive software solution designed to handle the entire pipeline of omics data processing. GenoCraft offers a unified platform featuring advanced bioinformatics tools, covering all aspects of omics data analysis. It encompasses a range of functionalities, such as normalization, quality control, differential analysis, network analysis, pathway analysis, and diverse visualization techniques. This software makes state-of-the-art omics data analysis more accessible to a wider range of users. With GenoCraft, researchers and data scientists have access to an array of cutting-edge bioinformatics tools under a user-friendly interface, making it a valuable resource for managing and analyzing large-scale omics data.
The API with an interactive web interface is publicly available at https://genocraft.stanford.edu/. We also release all the codes in https://github.com/futianfan/GenoCraft.
@@ 1 Introduction
During the past few years, the field of genomics has seen a remarkable and unprecedented expansion, fueled by an enormous volume of data. High-throughput sequencing platforms have provided unprecedented opportunities to delve into the genomic landscape, heralding a surge in omics data, including genomics, transcriptomics, proteomics, and metabolomics [${}^{+}$06, ${}^{+}$22]. However, while generating omics data has become increasingly accessible, analyzing these complex datasets remains a significant challenge to researchers and practitioners.
The handling and interpretation of the vast amounts of data generated by high-throughput technologies necessitate sophisticated computational tools [CLS16]. An intricate analysis pipeline is required to gain meaningful insights into these data, including quality control, preprocessing, alignment, variant calling, differential expression analysis, pathway analysis, etc [${}^{+}$05]. In order to carry out each of these steps, researchers need a specialized bioinformatics tool. Many researchers find these tools difficult to learn, as they require substantial computational resources [HLB18].
GenoCraft aims to fill this gap by providing a comprehensive, intuitive platform that facilitates seamless navigation through complex omics data analysis. This web-based software solution integrates a suite of state-of-the-art bioinformatics tools within a user-friendly interface, thereby democratizing access to advanced omics data analysis. GenoCraft is not merely a tool; it is a versatile companion designed to empower researchers, enabling them to harness the full potential of omics data in driving breakthroughs in their respective fields [${}^{+}$16].
@@ 2.1 Quality Control
  • Ensuring data integrity: By checking the data quality, researchers can ensure that the data used in their analysis is accurate and reliable. This step is critical to validate the conclusions drawn from the analysis [${}^{+}$21].
@@ 2.2 Normalization
Normalization is a preliminary step in the analysis of high-throughput omics data [${}^{+}$15]. GenoCraft employs a sophisticated algorithm for normalization to correct for technical variations such as different sequencing depths or RNA composition, ensuring comparability across samples [${}^{+}$15, ${}^{+}$22]. The specific normalization technique used is dependent on the nature of the data, with options including Total Count, RPKM (reads per kilobase of transcript per million reads mapped)/FPKM (fragments per kilobase of transcript per million fragments mapped), TPM (transcripts per million), and others for RNA-seq, or TMM (Trimmed Mean of M-values) for gene expression [${}^{+}$23, ZYS20].
In the CPM (Counts Per Million) normalization procedure, the raw read counts for each gene in a sample are transformed into proportions relative to the total number of mapped reads per sample. This transformation allows for comparisons of gene expression levels across different samples. The CPM normalization method follows the approach described by Robinson et al. in [${}^{+}$].
@@ 2.3 Differential Analysis
Identifying markers in genomics and other omics studies is of paramount importance as these markers often represent genes, proteins, or other molecules that exhibit significant differences between conditions, such as healthy and diseased subtypes [${}^{+}$21].
Differential analysis is a core component of GenoCraft’s analytical repertoire. The software utilizes robust statistical models to identify genes, transcripts, or proteins that exhibit significant differences in expression between conditions [${}^{+}$22].
The t-test, integral to the process of marker identification in GenoCraft, is leveraged to compare the mean values between two distinct groups to ascertain their statistical divergence [${}^{+}$09]. Within the GenoCraft system, a predefined threshold is established to filter and highlight genes that surpass this level, demonstrating statistically significant differences, thereby ensuring robust and reliable identification of key markers.
@@ 2.4 Network Analysis
GenoCraft incorporates differential network analysis, known as Differential Dependency Networks (DDN3.0) [${}^{+}$21], which is skilled at concurrently learning both common and rewired network structures. Network analysis identifies in a complex and often unknown overall molecular circuitry a network of differentially connected molecular entities. Researchers can investigate these networks to uncover modules of interconnected genes or proteins, identify central or “hub” nodes, and explore how signals propagate through the network [${}^{+}$12], and further help to provide a plausible interpretation of data, gain new insight of disease biology, and generate novel hypotheses for further validation and investigations.
@@ 2.5 Gene Set Enrichment Analysis
By mapping genes onto known pathways from databases such as KEGG (Kyoto Encyclopedia of Genes and Genomes) [KG00] or Reactome [${}^{+}$18], GenoCraft provides insights into the biological mechanisms underlying the observed changes in gene or protein expression [${}^{+}$05]. The software also employs statistical methods to identify pathways that are significantly enriched (with p-values provided), providing a clear direction for further investigation [${}^{+}$16].
@@ 2.6 Clustering
A clustering algorithm is a machine learning technique used to group similar data points together based on their inherent similarities or patterns [HW79]. It aims to partition a dataset into subsets or clusters, where data points within each cluster are more similar to each other than to those in other clusters [${}^{+}$17]. In the context of omics data, clustering techniques categorize genes, proteins, or samples into groups, such that entities within a cluster show higher similarity to each other based on the chosen metric and less similarity to entities in other clusters [${}^{+}$11].
@@ 3.1 Case Study: Bulk RNA-seq Analysis of GSE99611
GSE99611 (https://ftp.ncbi.nlm.nih.gov/geo/series/GSE99nnn/GSE99611/matrix/) is a bulk RNA-sequencing dataset available on the Gene Expression Omnibus (GEO) database. It contains gene expression information generated from a specific biological study [${}^{+}$10]. Our aim in this case study is to elucidate the underlying biological insights from this dataset by using a comprehensive RNA-seq analysis pipeline.
@@ 3.2 Case Study 2: Single-Cell Analysis of GSE69405
The dataset GSE69405 (https://ftp.ncbi.nlm.nih.gov/geo/series/GSE69nnn/GSE69405/matrix/) presents a rich resource for understanding the intricacies of lung adenocarcinoma at the single-cell level. This dataset leverages high-throughput sequencing to profile the transcriptomes of single cancer cells derived from patient lung adenocarcinoma xenografts (PDX) tumors [${}^{+}$95]. The overarching aim was to investigate the impact of intratumoral heterogeneity on anti-cancer drug responses.
@@ 3.3 Case Study 3: Protein
In this case study, we present a comprehensive pipeline for protein analysis using simulated data of 100 proteins and 50 samples. The pipeline encompasses various steps, including data preprocessing, quality control, differential expression analysis, functional enrichment analysis, and visualization. We demonstrate the effectiveness of the pipeline by applying it to the Genocraft application, showcasing its potential for protein research and discovery.
First, we performed quality control measures to identify proteins with low expression counts and excluded them from subsequent analyses. This step improved the robustness and reliability of our results. Then we preprocessed the data by handling missing values, and normalized expression data.
Following quality control, we conducted differential expression analysis to identify proteins showing significant differences in expression between samples or experimental conditions. T-tests were employed to detect these differentially expressed proteins, as shown in Figure 5.
Figure 5: Protein pathway analysis.
To gain further insights into the biological relevance of the differentially expressed proteins, we performed gene set enrichment analysis to identify enriched terms and pathways associated with the differentially expressed proteins. This step sheds light on the underlying biological processes and pathways implicated in the dataset, as shown in Figure 6.
Figure 6: Protein pathway analysis.
The algorithm of Genocraft also applied in the identification of markers in the Proteomic Architecture of Human Coronary and Aortic Atherosclerosis [${}^{+}$18], published in Circulation.
@@ 5 Discussion
GenoCraft, a state-of-the-art web-based platform, stands as a formidable powerhouse for high-throughput omics data analysis and visualization. By synergizing a diverse array of cutting-edge bioinformatics techniques, it has anchored its position as an indispensable resource for researchers delving deep into the intricate maze of omics data. Right from the initial steps of normalization and quality assurance to the more intricate facets of differential, network, and pathway analyses, GenoCraft delivers an exhaustive and meticulous data scrutiny [${}^{+}$19, ${}^{+}$23].
@@ 8 Contributions
Y.L., T.F, and Y.W. developed the Genocraft method and algorithm. T.F. and F.M. applied the Genocraft framework and performed data analyses. Y.L. and C.V.R. interpreted the data and wrote the manuscript. M.S. and C.L. designed the user interface and the front-end and back-end implementation. D.H and C.V.R. supervised the study.

### v2

@@
Yingzhou Lu1,2, Minjie Shen2, Ling Yue3, Chenhao Li4, Fan Meng2, Xiao Wang5,
David Herrington6, Yue Wang2, Yue Zhao7, Tianfan Fu3, Capucine Van Rechem1,‡
1Department of Pathology, Stanford University School of Medicine, Stanford, CA, 94305, USA
2 Bradley Department of Electrical and Computer Engineering, Virginia Tech, Blacksburg, VA, 24060, USA
3Department of Computer Science, Rensselaer Polytechnic Institute, Troy, NY, 12180, USA
4 Computer Engineering, University of Illinois Urbana-Champaign, Urbana, IL, 61820, USA
5Department of Computer Science and Engineering, University of Washington, 1410 NE Campus Parkway Seattle, WA 98195, USA
6Department of Internal Medicine, Wake Forest University, Winston-Salem, NC, 27109, USA
7 Department of Computer Science, University of Southern California, Los Angeles, CA, 90089, USA
‡Lead contact: cvrechem@stanford.edu
@@ 7 Department of Computer Science, University of Southern California, Los Angeles, CA, 90089, USA
The surge in high-throughput omics data has reshaped the landscape of biological research, underlining the need for powerful, user-friendly data analysis and interpretation tools. This paper presents GenoCraft, a web-based comprehensive software solution designed to handle the entire pipeline of omics data processing. GenoCraft offers a unified platform featuring advanced bioinformatics tools, covering all aspects of omics data analysis. It encompasses a range of functionalities, such as normalization, quality control, differential analysis, network analysis, pathway analysis, and diverse visualization techniques. This software makes state-of-the-art omics data analysis more accessible to a wider range of users. With GenoCraft, researchers and data scientists have access to an array of cutting-edge bioinformatics tools under a user-friendly interface, making it a valuable resource for managing and analyzing large-scale omics data. The API with an interactive web interface is publicly available at https://genocraft.stanford.edu/. We also release all the codes in https://github.com/futianfan/GenoCraft.
@@ 1 Introduction
During the past few years, the field of genomics has seen a remarkable and unprecedented expansion, fueled by an enormous volume of data. High-throughput sequencing platforms have provided unprecedented opportunities to delve into the genomic landscape, heralding a surge in omics data, including genomics, transcriptomics, proteomics, and metabolomics [REG+06, HFG+22]. However, while generating omics data has become increasingly accessible, analyzing these complex datasets remains a significant challenge to researchers and practitioners.
The handling and interpretation of the vast amounts of data generated by high-throughput technologies necessitate sophisticated computational tools [CLS16]. An intricate analysis pipeline is required to gain meaningful insights into these data, including quality control, preprocessing, alignment, variant calling, differential expression analysis, pathway analysis, etc [STM+05]. In order to carry out each of these steps, researchers need a specialized bioinformatics tool. Many researchers find these tools difficult to learn, as they require substantial computational resources [HLB18].
GenoCraft aims to fill this gap by providing a comprehensive, intuitive platform that facilitates seamless navigation through complex omics data analysis. This web-based software solution integrates a suite of state-of-the-art bioinformatics tools within a user-friendly interface, thereby democratizing access to advanced omics data analysis. GenoCraft is not merely a tool; it is a versatile companion designed to empower researchers, enabling them to harness the full potential of omics data in driving breakthroughs in their respective fields [IKK+16].
@@ 2.1 Quality Control
  • Ensuring data integrity: By checking the data quality, researchers can ensure that the data used in their analysis is accurate and reliable. This step is critical to validate the conclusions drawn from the analysis [YOW+21].
@@ 2.2 Normalization
Normalization is a preliminary step in the analysis of high-throughput omics data [DZZ+15]. GenoCraft employs a sophisticated algorithm for normalization to correct for technical variations such as different sequencing depths or RNA composition, ensuring comparability across samples [DZZ+15, WSD+22]. The specific normalization technique used is dependent on the nature of the data, with options including Total Count, RPKM (reads per kilobase of transcript per million reads mapped)/FPKM (fragments per kilobase of transcript per million fragments mapped), TPM (transcripts per million), and others for RNA-seq, or TMM (Trimmed Mean of M-values) for gene expression [DBP+23, ZYS20].
In the CPM (Counts Per Million) normalization procedure, the raw read counts for each gene in a sample are transformed into proportions relative to the total number of mapped reads per sample. This transformation allows for comparisons of gene expression levels across different samples. The CPM normalization method follows the approach described by Robinson et al. in [AMC+].
@@ 2.3 Differential Analysis
Identifying markers in genomics and other omics studies is of paramount importance as these markers often represent genes, proteins, or other molecules that exhibit significant differences between conditions, such as healthy and diseased subtypes [CLW+21].
Differential analysis is a core component of GenoCraft’s analytical repertoire. The software utilizes robust statistical models to identify genes, transcripts, or proteins that exhibit significant differences in expression between conditions [LWP+22].
The t-test, integral to the process of marker identification in GenoCraft, is leveraged to compare the mean values between two distinct groups to ascertain their statistical divergence [RSS+09]. Within the GenoCraft system, a predefined threshold is established to filter and highlight genes that surpass this level, demonstrating statistically significant differences, thereby ensuring robust and reliable identification of key markers.
@@ 2.4 Network Analysis
GenoCraft incorporates differential network analysis, known as Differential Dependency Networks (DDN3.0) [ZFL+21], which is skilled at concurrently learning both common and rewired network structures. Network analysis identifies in a complex and often unknown overall molecular circuitry a network of differentially connected molecular entities. Researchers can investigate these networks to uncover modules of interconnected genes or proteins, identify central or “hub” nodes, and explore how signals propagate through the network [MCK+12], and further help to provide a plausible interpretation of data, gain new insight of disease biology, and generate novel hypotheses for further validation and investigations.
@@ 2.5 Gene Set Enrichment Analysis
By mapping genes onto known pathways from databases such as KEGG (Kyoto Encyclopedia of Genes and Genomes) [KG00] or Reactome [FJM+18], GenoCraft provides insights into the biological mechanisms underlying the observed changes in gene or protein expression [STM+05]. The software also employs statistical methods to identify pathways that are significantly enriched (with p-values provided), providing a clear direction for further investigation [KJR+16].
@@ 2.6 Clustering
A clustering algorithm is a machine learning technique used to group similar data points together based on their inherent similarities or patterns [HW79]. It aims to partition a dataset into subsets or clusters, where data points within each cluster are more similar to each other than to those in other clusters [AGBM+17]. In the context of omics data, clustering techniques categorize genes, proteins, or samples into groups, such that entities within a cluster show higher similarity to each other based on the chosen metric and less similarity to entities in other clusters [PSM+11].
@@ 3.1 Case Study: Bulk RNA-seq Analysis of GSE99611
GSE99611 (https://ftp.ncbi.nlm.nih.gov/geo/series/GSE99nnn/GSE99611/matrix/) is a bulk RNA-sequencing dataset available on the Gene Expression Omnibus (GEO) database. It contains gene expression information generated from a specific biological study [BTW+10]. Our aim in this case study is to elucidate the underlying biological insights from this dataset by using a comprehensive RNA-seq analysis pipeline.
@@ 3.2 Case Study 2: Single-Cell Analysis of GSE69405
The dataset GSE69405 (https://ftp.ncbi.nlm.nih.gov/geo/series/GSE69nnn/GSE69405/matrix/) presents a rich resource for understanding the intricacies of lung adenocarcinoma at the single-cell level. This dataset leverages high-throughput sequencing to profile the transcriptomes of single cancer cells derived from patient lung adenocarcinoma xenografts (PDX) tumors [HSH+95]. The overarching aim was to investigate the impact of intratumoral heterogeneity on anti-cancer drug responses.
@@ 3.3 Case Study 3: Protein
Rhabdoid tumors, highly aggressive childhood cancers, are driven by the loss of the mSWI/SNF (mammalian SWItch/Sucrose Non-Fermentable) subunit SMARCB1. These tumors, which can develop in the brain, kidneys, or soft tissues, exhibit specific sensitivity to the translation inhibitor homoharringtonin (HHT).
In this computational biology-driven analysis, we utilized GenoCraft to perform a comprehensive examination of protein expression data from rhabdoid tumor cells, focusing on identifying significant proteins and their associated pathways. The protein dataset was first subjected to quality control and normalization processes to ensure the reliability of the results.
The differential expression analysis revealed a set of 24 proteins with statistically significant changes between the conditions under study. These proteins were identified based on an adjusted p-value threshold of 0.02 and a log2 fold change (LogFC) greater than 1.5 for upregulated proteins. The volcano plot generated from this analysis provided a clear visual representation of the significant proteins, with those meeting the criteria for differential expression highlighted in red. Notably, proteins associated with chromatin remodeling, particularly those within the SWI/SNF complex, emerged as key players in the dataset, suggesting their involvement in critical cellular processes.
Further pathway enrichment analysis provided additional insights into the biological functions of the significant proteins. The top 20 pathways, ranked by -log10(Adjusted p-value), included the SWI/SNF complex, ATPase complex, and pathways related to chromatin remodeling. These findings underscore the potential role of these pathways in driving the observed protein expression changes. By integrating differential expression analysis with pathway enrichment from GO, this case study highlights the utility of GenoCraft in uncovering biologically relevant insights that can inform future research directions and therapeutic development.
The results of this analysis not only enhance our understanding of the molecular mechanisms underlying the studied conditions but also demonstrate the effectiveness of GenoCraft as a comprehensive tool for high-throughput omics data analysis. The identification of key proteins and pathways provides a strong foundation for further exploration, potentially leading to the discovery of novel biomarkers or therapeutic targets.
Figure 5: significant protein markers
Figure 6: Top 20 GO Pathways
@@ 5 Discussion
GenoCraft, a state-of-the-art web-based platform, stands as a formidable powerhouse for high-throughput omics data analysis and visualization. By synergizing a diverse array of cutting-edge bioinformatics techniques, it has anchored its position as an indispensable resource for researchers delving deep into the intricate maze of omics data. Right from the initial steps of normalization and quality assurance to the more intricate facets of differential, network, and pathway analyses, GenoCraft delivers an exhaustive and meticulous data scrutiny [LCH+19, WFD+23].
@@ 8 Contributions
Y.L., T.F, and Y.W. developed the Genocraft method and algorithm. T.F. and F.M. applied the Genocraft framework and performed data analyses. Y.L. and C.V.R. interpreted the data and wrote the manuscript. M.S., L.Y. and T.F. designed the user interface and the front-end and back-end implementation. D.H and C.V.R. supervised the study.